OpenStudy (anonymous):

i have all the data i just need someone to help explain it because this question is confusing

OpenStudy (anonymous):

Whats your questionn?

OpenStudy (anonymous):

What is the average common ratio between the successive height values of ball 1? Ball 2? Ball 3? Experimental errors may cause common ratios to have some variances within the data for one ball. Use the average common ratio.

OpenStudy (anonymous):

thats my data

OpenStudy (anonymous):

ohhh im not sure about thiss.

OpenStudy (anonymous):

For the first one, you need to decide "what do I have to do to the 3 to get to the 2.4? Then what do I have to do to the 2.4 to get to the 1.9?" What I did to come up with answer was divide the 3 by 2.4. I came up with 1.25. So I decided to see if, when I divide each term by 1.25, will I get the next term in line? I did. 3/1.25 = 2.4. 2.4/1.25= 1.9. 1.9/1.25 = 1.5. So the common rati 1.58 is ball 2's common ratio. I found that the same way. See, the common ratio is only a reflection of the change in the data. The change from one point to the next is to divide it by 1.58. Here's from you problem above: Experimental errors may cause common ratios to have some variances within the data for one ball. This must be that ball! The average common ratio would be to add the 2 ratios you found and divide them by 2. That's the average common ratio.

OpenStudy (anonymous):

listen to himm^ lol

OpenStudy (anonymous):

okay next one and thank you @ameenbriscoe13

OpenStudy (anonymous):

What is the height of each ball on the fifth bounce (i.e., Height 6)? Use the geometric sequence formula, an = a1rn – 1 and show your work.

OpenStudy (anonymous):

\[a _{n}=a _{1}r ^{n-1}\]

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

Well ball bounce 1 I'm guessing is